OpenStudy (anonymous):

Use the comparison theorem to determine whether the integral of sec^2x / xsqrt(x) dx from 0 to 1 is convergent or divergent

ganeshie8 (ganeshie8):

Hint : since \(|\sec x| \ge 1\), we have : \(\dfrac{\sec^2x}{x\sqrt{x}} \ge\dfrac{1}{x\sqrt{x}} \)

OpenStudy (anonymous):

I understood that, but how do I prove it, and I want to understand how you can determine why something diverges or is convergent.

ganeshie8 (ganeshie8):

good question, let me ask you a quick question, look at below sum and tell me if you think it converges : \[ 1+1+1+1+\cdots \]

OpenStudy (anonymous):

yes because it is greater than one

ganeshie8 (ganeshie8):

so are you saying that sum converges ?

OpenStudy (anonymous):

that's what I thought, am I wrong?

ganeshie8 (ganeshie8):

yes, loosely speaking, "converging sum" means that the sum approaches some value as you increase the number of terms that you add

ganeshie8 (ganeshie8):

\[ 1+1+1+1+\cdots \] look at above sum again, does it seem like approaching any finite number ?

ganeshie8 (ganeshie8):

maybe take the partial sums and convince yourself that the sum is NOT approaching any fixed number : 1+1 = ? 1+1+1 = ? 1+1+1+1 = ? 1+1+1+1+1 = ?

OpenStudy (anonymous):

I get it now, the sum will be approaching infinity and so therefore it has to be diverging

ganeshie8 (ganeshie8):

Exactly! You got it mostly! but thats one way to look at it.

ganeshie8 (ganeshie8):

look at below sum and tell me if it converges or diverges : \[1-1+1-1+1-1+\cdots \]

OpenStudy (anonymous):

i would think it still diverges because it is not stable

ganeshie8 (ganeshie8):

Yep! now you know what the words, "converge" or "diverge" mean exactly. Lets try one more last example

ganeshie8 (ganeshie8):

Look at below sum, does it converge ? \[32+ 16+ 8+4+2+1+\dfrac{1}{2}+\dfrac{1}{4}+\cdots\]

OpenStudy (anonymous):

no because regardless of the value it will still be constantly increasing and so it is not a standard limit

ganeshie8 (ganeshie8):

think again, here is a hint : geometric series

OpenStudy (anonymous):

I don't understand, I do notice that it is all going down at a ratio of 1/2

ganeshie8 (ganeshie8):

Yes, remember the partial sum formula for geometric series ?

ganeshie8 (ganeshie8):

if not, don't bother... we don't really need to know about geometric series as of now...

OpenStudy (anonymous):

I dont think I was taught that

ganeshie8 (ganeshie8):

that makes sense, okay we're done with examples. lets get back to the original problem

ganeshie8 (ganeshie8):

so, with integerals also, we define convergence and divergence similalrly : if the integral evaluates to a fixed number, we say the integral "converges" otherwise, we say the integral diverges

OpenStudy (anonymous):

I did the steps and it ended up being -2, but I know that it is supposed to diverge

ganeshie8 (ganeshie8):

what has ended up being -2 ?

OpenStudy (anonymous):

I'll write the steps that I took

ganeshie8 (ganeshie8):

Okay

OpenStudy (anonymous):

\[\lim_{t \rightarrow 0} \int\limits_{t}^{1} 1/(x \sqrt{x}) \] \[\int\limits_{t}^{1} x^-3/2\] \[\lim_{t \rightarrow 0} \left( -2x^(-\frac{ 1 }{ 2 } )\right)\] \[\lim_{t \rightarrow 0} \left[ -2 - (-2t^(-1/2) \right]\] = -2 when t goes to 0

ganeshie8 (ganeshie8):

Looks very good till the last line there is a huge mistake in the last line when you evaluate the limit

ganeshie8 (ganeshie8):

\(\lim\limits_{t \rightarrow 0} \left[ -2 - (-2t^{-1/2} \right]\) is same as \(\lim\limits_{t \rightarrow 0} \left[ -2 + \dfrac{2}{\sqrt{t}} \right]\) yes ?

OpenStudy (anonymous):

i thought because it went to 0 it got eliminated

ganeshie8 (ganeshie8):

Now if u evaluate the limit, you would see that you get infinity so the integral doesn't converge.

ganeshie8 (ganeshie8):

look at below limit : \(\lim\limits_{t\to 0} \dfrac{1}{t}\) what does it evaluate to ?

ganeshie8 (ganeshie8):

1/0.1 = ? 1/0.01 = ? 1/0.00001 = ?

OpenStudy (anonymous):

infinity >< I messed up, I was thinking like 0/1 = 0

ganeshie8 (ganeshie8):

Happens... Okay, so now we know that the integral \(\int\limits_0^1 \dfrac{1}{x\sqrt{x}}\,dx\) doesn't converge.

ganeshie8 (ganeshie8):

what does that tell us about the integral \(\int\limits_0^1 \dfrac{\sec^2x}{x\sqrt{x}}\,dx\) ?

OpenStudy (anonymous):

that it diverges also based on f(x) > g(x) and there is the concept that if g(x) diverges then so does f(x)

ganeshie8 (ganeshie8):

Looks good! when free just take a quick look http://tutorial.math.lamar.edu/Classes/CalcII/ImproperIntegralsCompTest.aspx

OpenStudy (anonymous):

Thank you so much for your help, my calculus 2 exam is like in 8 hours. I really appreciate it.

ganeshie8 (ganeshie8):

np, You seem to be mostly ready with the material ! good luck !